OpenStudy (anonymous):

A 4-sided shape has interior angles of 71 degrees 200 degrees 77 degrees and what is the other degree? Explainnn

OpenStudy (anonymous):

Sum of interior angles: sum = 180(n-2) where n is the number of sides

OpenStudy (anonymous):

Um Can you Explain more]

OpenStudy (mathstudent55):

For a polygon of n sides, the formula (n - 2)180 gives the sum of the measures of the interior angles. For example, a triangle is a polygon with 3 sides. What is the sum of the measures of the interior angles of a triangle? Here, n = 3, so (3 - 2)180 = 1(180) = 180 In the problem you posted, you have a 4-sided polygon. You know the measures of 3 of its 4 angles. Since you can calculate, using the above formula, the sum of the measures of all 4 angles, you can figure out the measure of the missing angle.